# Mohanpur, West Bengal 741246, India

### v1

@@ 1 Introduction
The paper is organized as follows: In Section 2, we give an overview of inflationary dynamics, with particular emphasis on the $\alpha$-attractor models of inflation. We describe the T-Model and Polynomial Attractor model in detail, which will be our primary focus in this paper. In Section 3 we describe in detail the nature of the gravitational wave spectrum from the local and global cosmic strings. In Section 4, we explore a scenario of cosmic string formation during inflation. We introduced the turning-point frequency within the framework of the VOS model of cosmic strings and derived an explicit expression for it. We also show the impact of T-Model inflation on the gravitational wave spectra of the local and global cosmic strings. In Section 5, we discuss our results, and provide a summary and conclusion of our work in Section 6.
@@ 2 Inflationary Models and Predictions
The current observational constraints on $n_{s}$, $n_{s}^{\prime}$ and $n_{s}^{\prime\prime}$ are given by Planck 2018 Planck:2018vyg : $n_{s}=0.964\pm 0.004$ (at $68$% C.L.), $n_{s}^{\prime}=0.001\pm 0.009$ and $n_{s}^{\prime\prime}=0.009\pm 0.012.$. The value of the tensor-to-scalar ratio, $r$, is constrained by the Planck, WMAP, and BICEP/Keck Observations BICEP:2021xfz : $r_{0.05}<0.036$ (at 95% C.L.). All these observations were made at the pivot scale, $k_{\star}=0.05\mathrm{Mpc}^{-1}$. Computing the slow-roll parameters at a fiducial value of the e-fold, $N_{e}$, defined as the number of e-folds between the horizon exit of the pivot CMB mode and the end of inflation, allows consistent approximations for $n_{s}$ and $r$ in cases when the slow roll approximation is valid. $N_{e}$ is defined as:
[…]
@@
where, $\rho_{*}=V(\phi_{*})$ is the energy density of the Universe when the pivot scale exits the horizon, $\rho_{end}=V(\phi_{end})$ is the energy density at the end of inflation and $\rho_{reh}=(\pi^{2}/30)g_{*}T_{reh}^{4}$ is the energy density at the reheating time and $\omega_{reh}$ is the effective equation-of-state parameter from the end of inflation until reheating. The effective number of massless degrees of freedom during the reheating epoch is taken to be $g_{*}=106.75$. In our analysis, we will consider instant reheating where $\omega_{reh}=1/3$.
@@ 2.1 $\alpha$-attractors
Figure 1: Potential profiles of the E-model $\alpha$-attractors given by Eq. (19), illustrating the dependence on the parameters $\alpha$ and $n$. (a) The potential $V(\phi)/V_{0}$ is plotted as a function of $\phi/M_{Pl}$ for fixed $n=1$ and varying $\alpha$. (b) The potential is shown for fixed $\alpha=1$ and varying $n$. These plots highlight the sensitivity of the inflationary potential’s shape to changes in $\alpha$ and $n$ the key parameters in the $\alpha$-attractor framework.
@@ 2.1.2 Polynomial $\alpha$-attractor
3 Gravitation Waves from Cosmic Strings: short review
[…]
3.1 The gravitational-wave spectrum from Local Cosmic Strings
$\mu=2\pi n\eta^{2},$ (33)
$P_{\mathrm{GW}}=\Gamma G\mu^{2},$ (35)
where
×$\times\left[\frac{a(t_{i})}{a(\tilde{t})}\right]^{3}\cdot\mathcal{P}_{\mathrm{loop}}(\alpha)\cdot\frac{dt_{i}}{df}\cdot\frac{dn_{\mathrm{loop}}}{dt_{i}}$
[…]
Figure 5 shows the full GW spectrum  [^2] We have used the publicly available code https://github.com/moinulrahat/CosmicStringGW Fu:2024rsm for plotting GW spectrum. due to local CS computed using Eq. (40) assuming a $\Lambda$CDM universe and the strings were formed after inflation ends. The spectrum can be broadly classified into a flat high-frequency regime and the radiation-to-matter transition at the low-frequency regime. Higher and higher frequencies correspond to earlier and earlier in time scales, i.e., it corresponds to a radiation-dominated era. The local CS loops that were produced during the radiation-dominated epoch and begun GW emission have a nearly flat GW spectrum, with the GW amplitude being expressed as,
$\tilde{t_{M}}=\frac{\alpha}{2\Gamma G\mu}t_{i}.$ (44)
3.2 The gravitational-wave spectrum from Global Cosmic Strings
$T_{form}^{g}\simeq\eta.$ (46)
where,
[…]
where,
$C_{\mathrm{eff}}^{\mathrm{g}}=\tilde{c}\bar{v}/\xi^{3},$ (51)
$\implies\quad\eta^{global}\lesssim 1.4\times 10^{15}GeV,$ (54)
where $t_{CMB}=374$ kyr.
A more detailed description of local and global strings can be found in Refs. Gouttenoire:2019kij ; Chang:2021afa ; Vilenkin:2000jqa
3.3 Gravitational Wave Detectors
In figures 5 and 6 we have depicted the present day and future planned GW experimental sensitivity reaches of the following categories:
@@ 4 Formation of Cosmic Strings during Inflation
Inflation smoothens out any initial inhomogeneities and drives the universe towards a nearly flat geometry. However, if inflation ends through a phase transition, as in many hybrid inflation models, it can lead to the formation of topological defects, including cosmic strings. Vilenkin:2000jqa
@@ 4.2 Turning-Point Frequency
An important observable in our analysis will be the turning-point frequency $f_{\Delta}$, above which the GW spectrum deviates from the one obtained in standard cosmology (Fig. 5). At higher frequencies, both the standard and the non-standard scenario show a characteristic $\Omega_{GW}h^{2}\sim f^{-1}$ behaviour, which happens to be a consequence of the lack of production of the loops that contributes dominantly in this frequency range. But in the non-standard scenario that we are going to discuss ahead, we will find that, GW spectra from cosmic strings that have experienced some e-foldings of inflation, shifts this $\Omega_{GW}h^{2}\sim f^{-1}$ behaviour to lower frequencies. Guedes:2018afo . Apart from understanding the non-standard scenario due to the impact of inflation that we will describe in this paper, another important non-standard scenario would be the presence of an early matter dominated era, see Refs. Cui:2017ufi ; Ghoshal:2023sfa ; Datta:2025yow for detailed analyses. The only way to distinguish between them is to estimate the inflation affected cosmic string GW spectrum (as the one in Fig.7) considering at least $10^{4}$ $k$-modes (which is beyond the scope of this paper), which shows a transition from a $f^{-1/3}$ to $f^{-1}$ in the case of inflation Gouttenoire:2019kij . After the end of inflation, one must wait for the correlation length to re-enter the horizon in order to reach the scaling regime again. The duration of the transient regime receives an enhancement factor $\exp({N_{e}})$. The turning-point frequency for local cosmic strings following Nambu-Goto dynamics read as Gouttenoire:2019kij ; Cui:2018rwi ; Ghoshal:2023sfa ; Datta:2025yow :
@@ 4.3 Impact of a T-Model Inflation on GW spectrum due to local and global Cosmic String Loops
In figures 7, 8, 9 and 10 we have our modified GW spectra from local CS that were formed during inflation and has experienced a certain e-fold of inflation. From the plots, we can determine the points from which the modified spectra deviate from the ‘standard’ spectra, depending on how many e-folds of inflation it has suffered. This is where the turning-point frequency (given by Eq. (75)) lies. If our CS has suffered lesser e-folds of inflation, then its corresponding GW spectrum (formed due to the emission of GWs after the string network re-enters the horizon) deviates from the ‘standard’ behaviour at higher frequencies. Similarly, if it has suffered a greater amount of inflation, the spectra start to deviate right away from the low-frequency regime. Also, the more inflation the strings suffer, the narrower the flat plateau region becomes in the GW spectrum emitted by these strings. From the figures mentioned above, one can also point out that as the energy scale of inflation $E_{inf}$ (controlled by $r$) decreases, the maximum permissible e-fold of inflation $N_{CS}$ that the string undergoes also decreases. The $f^{-1}$ behaviour at higher frequencies of the non-standard GW spectrum for both local and global CS is a result of the late entry time of the string network in the horizon, which delays the time of loop production. Once the rate of loop production by both the standard and non-standard networks becomes equal as they reach the linear scaling regime, their low-frequency spectra coincide Guedes:2018afo . We will emphasize more about the bounds on $N_{CS}$ later on in section 5.
@@ 5.1 Local Cosmic Strings
Figures 12, 13, 14 and 15 provide crucial information connecting inflationary physics to observable GW signatures. It captures the range of inflationary parameters ($n_{s}$, $r$) that enables cosmic strings (formed during inflation) to produce observable GWs which can be detected at the present and future observatories. Consider the top-left plot in Fig. 15, which considers the Polynomial attractor inflation model with $n=1$ (using Eq. (27):
@@ 5.2 Global Cosmic Strings
$f_{\Delta}$ is in Hertz and $N_{CS}$ is the number of e-folds of inflation that the cosmic string has suffered. Interestingly enough, the complementarity relation for global strings is independent of the string tension. The shaded region in Figs. 16 and 17, just like in the case for local strings, correspond to those combinations of $n_{s}$ and $r$ that can produce detectable GW signals from Global CS (formed during inflation) at the current and future detectors. This mapping from the ($n_{s}$, $r$)-plane to the observable GW signal solely depends on the dynamics of inflation and the e-folds experienced by the strings ($N_{CS}$), making it a more effective probe of inflationary physics, as it depends on fewer parameters. Consider the top left plot from Fig. 16 which is based on the Polynomial attractor inflation model (see Eq. (27)) with $n=1$. It maps the $n_{s}-r$ parameter space to the observable GW signature at the SKA, assuming the VEV $\eta^{global}=10^{15}$ GeV. The gradient of the shaded region tells us that the SKA can probe $N_{CS}$ in the range $\sim 36$ to $46$. Similarly, from the other plots from the same figure, we can infer that LISA can probe $N_{CS}$ in the range $\sim 24$ to $37$, DECIGO/BBO in the range $\sim 20$ to $32$ and ET in the range $\sim 14$ to $24$. However, the parameter space of ET in our analysis is far beyond the reach of the current and future CMB experiments and hence can not be detected by them. But, if we can see observable GW signatures or in other words if we observe the turning point frequency of global strings with $\eta^{global}=10^{15}$ GeV at the ET, it would allow us to put new constraints on the $n_{s}-r$ parameter space.
@@ 6 Discussion and Conclusion
  • The T-Model and Polynomial attractor models of inflation have similar $n_{s}-r$ predictions in the CMB detectors, especially in the large $N_{e}$ and small $\alpha$ limit. However, we showed that this degeneracy can be broken if cosmic strings are formed during inflation and suffer a few e-folds of inflation. We use this information to create our inflation model and determine the estimated GW spectrum from the cosmic strings (see Eqs. (82) and (85) and Figs. 7 to 11.
@@
  • Eq. (82) gives the complementary relation for local cosmic strings which depends on the turning point frequency $f_{\Delta}$, the number of e-folds of inflation the local cosmic strings suffer $N_{CS}$ and the string tension $G\mu$ determined by the vacuum expectation value of the scalar field $\eta^{local}$. Figs. 12-15 show the complementarity plots which form a bridge between inflationary observables and GW signatures. In addition to putting bounds on the $N_{CS}$ values, one can also use this method to impose constraints on the permissible $n_{s}-r$ values that the inflation model can take. For instance, LISA can probe $N_{CS}\sim 34-47$ irrespective of the value of the local string tension. For the Polynomial Attractor model, consider the complementary plots in Fig.15: it allows us to constrain the spectral index $n_{s}$ within the range $0.962\lesssim n_{s}\lesssim 0.972$ for $n=1$, $0.956\lesssim n_{s}\lesssim 0.968$ for $n=2$, $0.954\lesssim n_{s}\lesssim 0.965$ for $n=3$ and $0.963\lesssim n_{s}\lesssim 0.964$ for $n=4$.
  • Eq. (85) gives the complementary relation for global cosmic strings and depends only on the turning point frequency $f_{\Delta}$ and the number of e-folds of inflation the local cosmic strings experience $N_{CS}$. Figs. 16 and 17 show the complementarity plots due to global cosmic strings. Besides putting constraints on the values of $N_{CS}$ for the different GW detectors, using this technique (just like the case of local cosmic strings) we can also put constraints and probe regions of the $n_{s}-r$ parameter space that are within the reach of the current and future CMB experiments. As an example the top right plot in Fig. 16 which shows the complementarity for global cosmic strings at the LISA, constrains the spectral index $n_{s}$ in the range $0.960\lesssim n_{s}\lesssim 0.963$ for the Polynomial Attractor model with $n=1$. A part of the aforementioned range of spectral index (corresponding to $N_{CS}\sim 30-37$) is consistent with Planck and lies well within the reach of next generation CMB experiments like CMB-S4 and LiteBIRD.

### v2

@@ 1 Introduction
The paper is organized as follows: In Section 2, we give an overview of inflationary dynamics, with particular emphasis on the $\alpha$-attractor models of inflation. We describe the T-Model and Polynomial Attractor model in detail, which will be our primary focus in this paper. In Section 3, we explore a scenario of cosmic string formation during inflation. We introduced the turning-point frequency within the framework of the VOS model of cosmic strings and derived an explicit expression for it. We also show the impact of T-Model inflation on the gravitational wave spectra of the local and global cosmic strings. In Section 4, we discuss our results, and provide a summary and conclusion of our work in Section 5. In Appendix 6 we describe in detail the nature of the gravitational wave spectrum from the local and global cosmic strings.
@@ 2 Inflationary Models and Predictions
The current observational constraints on $n_{s}$, $n_{s}^{\prime}$ and $n_{s}^{\prime\prime}$ are given by Planck 2018 Planck:2018vyg: $n_{s}=0.964\pm 0.004$ (at 6868% C.L.), $n_{s}^{\prime}=0.001\pm 0.009$ and $n_{s}^{\prime\prime}=0.009\pm 0.012.$. The value of the tensor-to-scalar ratio, $r$, is constrained by the Planck, WMAP, and BICEP/Keck Observations BICEP:2021xfz: $r_{0.05}<0.036$ (at 95% C.L.). All these observations were made at the pivot scale, $k_{\star}=0.05\mathrm{Mpc}^{-1}$. Computing the slow-roll parameters at a fiducial value of the e-fold, $N_{e}$, defined as the number of e-folds between the horizon exit of the pivot CMB mode and the end of inflation, allows consistent approximations for $n_{s}$ and $r$ in cases when the slow roll approximation is valid. $N_{e}$ is defined as:
$N_{e}=\frac{1}{M_{Pl}^{2}}\int_{\phi_{\mathrm{end}}}^{\phi}\frac{V}{V^{\prime}}d\phi=\frac{1}{M_{Pl}}\int_{\phi}^{\phi_{\mathrm{end}}}\frac{|d\phi|}{\sqrt{2\epsilon_{V}}}$ (12)
@@
where, $\rho_{*}=V(\phi_{*})$ is the energy density of the Universe when the pivot scale exits the horizon, $\rho_{end}=V(\phi_{end})$ is the energy density at the end of inflation and $\rho_{reh}=(\pi^{2}/30)g_{*}T_{reh}^{4}$ is the energy density at the reheating time and $\omega_{reh}$ is the effective equation-of-state parameter from the end of inflation until reheating. The effective number of massless degrees of freedom during the reheating epoch is taken to be $g_{*}=106.75$. In our analysis, we will consider instant reheating where $\omega_{reh}=1/3$ [^2] Instantaneous reheating is where all the energy of the potential is converted into radiation very quickly. A simple condition can be understood for the assumption behind instantaneous reheating in terms of the parameters that $\tau_{\phi}\leq t_{p}$, where $t_{p}$ is the age of the universe during the reheating, leads to $\varphi_{\rm end}\geq O(10^{18})$ GeV or $\rho_{\rm end}\geq O(10^{13})$ GeV, which is typical energy scales of inflation at the end, for $m_{\phi}=0.1\varphi_{\rm end}$. Values of $\varphi_{\rm end}$ smaller than these should be analyzed in more detail, and may deviate from instantaneous reheating which is beyond the scope of the present paper. Eqn (2.14) still holds to be true a value different from $w_{\rm reh}=1/3$ will change the no. of e-folds which predict eventually the exact GW spectrum shape since it depends on total $N_{\rm efolds}$ suffered by the cosmic string networks. In cases where reheating is not instantaneous $w_{\rm reh}$ represents the average value of the barotropic parameter $w$ over the entire period of reheating. It is needless to mention that the reheating sector can be completely separate from the inflaton, for instance, in various mechanisms of reheating where a separate species is responsible for reheating. like for examples via mechanisms like gravitational particle production also known as gravitational reheating Ford:1986sy; Chun:2009yu or instant preheating Felder:1998vq; Dimopoulos:2017tud, curvaton reheating Feng:2002nb; BuenoSanchez:2007jxm, Ricci reheating Dimopoulos:2018wfg; Opferkuch:2019zbd; Bettoni:2021zhq; Laverda:2023uqv; Figueroa:2024asq, reheating via evaporation of primordial black holes Dalianis:2021dbs; Bhaumik:2022pil; Bhaumik:2022zdd; Ghoshal:2023fno..
@@ 2.1 $\alpha$-attractors
Figure 1: Potential profiles of the E-model $\alpha$-attractors given by Eq. (19), illustrating the dependence on the parameters $\alpha$ and $n$. (a) The potential $V(\phi)/V_{0}$ is plotted as a function of $\phi/M_{Pl}$ for fixed $n=1$ and varying $\alpha$. (b) The potential is shown for fixed $\alpha=1$ and varying $n$. These plots highlight the sensitivity of the inflationary potential’s shape to changes in $\alpha$ and $n$ - the key parameters in the $\alpha$-attractor framework.
@@ 3 Formation of Cosmic Strings during Inflation
Inflation smoothens out any initial inhomogeneities and drives the universe towards a nearly flat geometry. However, if inflation ends through a phase transition, as in many hybrid inflation models, it can lead to the formation of topological defects, including cosmic strings Vilenkin:2000jqa.
@@ 3.2 Turning-Point Frequency
We refer the reader to the Appendix for a short review of generation of Gravitational Waves from local and global string network. An important observable in our analysis will be the turning-point frequency $f_{\Delta}$, above which the GW spectrum deviates from the one obtained in standard cosmology (Fig. 16 in Appendix 6). At higher frequencies, both the standard and the non-standard scenario show a characteristic $\Omega_{GW}h^{2}\sim f^{-1}$ behaviour, which happens to be a consequence of the lack of production of the loops that contributes dominantly in this frequency range. But in the non-standard scenario that we are going to discuss ahead, we will find that, GW spectra from cosmic strings that have experienced some e-foldings of inflation, shifts this $\Omega_{GW}h^{2}\sim f^{-1}$ behaviour to lower frequencies. Guedes:2018afo. Apart from understanding the non-standard scenario due to the impact of inflation that we will describe in this paper, another important non-standard scenario would be the presence of an early matter dominated era, see Refs. Cui:2017ufi; Ghoshal:2023sfa; Datta:2025yow for detailed analyses. The only way to distinguish between them is to estimate the inflation affected cosmic string GW spectrum (as the one in Fig.5) considering at least $10^{4}$ $k$-modes (which is beyond the scope of this paper), which shows a transition from a $f^{-1/3}$ to $f^{-1}$ in the case of inflation Gouttenoire:2019kij. After the end of inflation, one must wait for the correlation length to re-enter the horizon in order to reach the scaling regime again. The duration of the transient regime receives an enhancement factor $\exp({N_{e}})$. The turning-point frequency for local cosmic strings following Nambu-Goto dynamics read as Gouttenoire:2019kij; Cui:2018rwi; Ghoshal:2023sfa; Datta:2025yow:
@@ 3.3 Impact of a T-Model Inflation on GW spectrum due to local and global Cosmic String Loops
In figures 5, 6, 7 and 8 we have our modified GW spectra from local CS that were formed during inflation and has experienced a certain e-fold of inflation. From the plots, we can determine the points from which the modified spectra deviate from the ‘standard’ spectra, depending on how many e-folds of inflation it has suffered. This is where the turning-point frequency (given by Eq. (49)) lies. If our CS has suffered lesser e-folds of inflation, then its corresponding GW spectrum (formed due to the emission of GWs after the string network re-enters the horizon) deviates from the ‘standard’ behaviour at higher frequencies. Similarly, if it has suffered a greater amount of inflation, the spectra start to deviate right away from the low-frequency regime. Also, the more inflation the strings suffer, the narrower the flat plateau region becomes in the GW spectrum emitted by these strings. From the figures mentioned above, one can also point out that as the energy scale of inflation $E_{inf}$ (controlled by $r$) decreases, the maximum permissible e-fold of inflation $N_{CS}$ that the string undergoes also decreases. The $f^{-1}$ behaviour at higher frequencies of the non-standard GW spectrum for both local and global CS is a result of the late entry time of the string network in the horizon, which delays the time of loop production. Once the rate of loop production by both the standard and non-standard networks becomes equal as they reach the linear scaling regime, their low-frequency spectra coincide Guedes:2018afo. We will emphasize more about the bounds on $N_{CS}$ later on in section 4.
@@ 4.1 Local Cosmic Strings
Figures 10, 11, 12 and 13 provide crucial information connecting inflationary physics to observable GW signatures. It captures the range of inflationary parameters ($n_{s}$, $r$) that enables cosmic strings (formed during inflation) to produce observable GWs which can be detected at the present and future observatories. Consider the top-left plot in Fig. 13, which considers the Polynomial attractor inflation model with $n=1$ (using Eq. (27):
@@ 4.2 Global Cosmic Strings
$f_{\Delta}$ is in Hertz and $N_{CS}$ is the number of e-folds of inflation that the cosmic string has suffered. Interestingly enough, the complementarity relation for global strings is independent of the string tension. The shaded region in Figs. 14 and 15, just like in the case for local strings, correspond to those combinations of $n_{s}$ and $r$ that can produce detectable GW signals from Global CS (formed during inflation) at the current and future detectors. This mapping from the ($n_{s}$, $r$)-plane to the observable GW signal solely depends on the dynamics of inflation and the e-folds experienced by the strings ($N_{CS}$), making it a more effective probe of inflationary physics, as it depends on fewer parameters. Consider the top left plot from Fig. 14 which is based on the Polynomial attractor inflation model (see Eq. (27)) with $n=1$. It maps the $n_{s}-r$ parameter space to the observable GW signature at the SKA, assuming the VEV $\eta^{global}=10^{15}$ GeV. The gradient of the shaded region tells us that the SKA can probe $N_{CS}$ in the range $\sim 36$ to $46$. Similarly, from the other plots from the same figure, we can infer that LISA can probe $N_{CS}$ in the range $\sim 24$ to $37$, DECIGO/BBO in the range $\sim 20$ to $32$ and ET in the range $\sim 14$ to $24$. However, the parameter space of ET in our analysis is far beyond the reach of the current and future CMB experiments and hence can not be detected by them. But, if we can see observable GW signatures or in other words if we observe the turning point frequency of global strings with $\eta^{global}=10^{15}$ GeV at the ET, it would allow us to put new constraints on the $n_{s}-r$ parameter space.
@@ 5 Discussion and Conclusion
  • The T-Model and Polynomial attractor models of inflation have similar $n_{s}-r$ predictions in the CMB detectors, especially in the large $N_{e}$ and small $\alpha$ limit. However, we showed that this degeneracy can be broken if cosmic strings are formed during inflation and suffer a few e-folds of inflation. We use this information to create our inflation model and determine the estimated GW spectrum from the cosmic strings (see Eqs. (56) and (59) and Figs. 5 to 9.
@@
  • Eq. (56) gives the complementary relation for local cosmic strings which depends on the turning point frequency $f_{\Delta}$, the number of e-folds of inflation the local cosmic strings suffer $N_{CS}$ and the string tension $G\mu$ determined by the vacuum expectation value of the scalar field $\eta^{local}$. Figs. 10-13 show the complementarity plots which form a bridge between inflationary observables and GW signatures. In addition to putting bounds on the $N_{CS}$ values, one can also use this method to impose constraints on the permissible $n_{s}-r$ values that the inflation model can take. For instance, LISA can probe $N_{CS}\sim 34-47$ irrespective of the value of the local string tension. For the Polynomial Attractor model, consider the complementary plots in Fig.13: it allows us to constrain the spectral index $n_{s}$ within the range $0.962\lesssim n_{s}\lesssim 0.972$ for $n=1$, $0.956\lesssim n_{s}\lesssim 0.968$ for $n=2$, $0.954\lesssim n_{s}\lesssim 0.965$ for $n=3$ and $0.963\lesssim n_{s}\lesssim 0.964$ for $n=4$.
  • Eq. (59) gives the complementary relation for global cosmic strings and depends only on the turning point frequency $f_{\Delta}$ and the number of e-folds of inflation the local cosmic strings experience $N_{CS}$. Figs. 14 and 15 show the complementarity plots due to global cosmic strings. Besides putting constraints on the values of $N_{CS}$ for the different GW detectors, using this technique (just like the case of local cosmic strings) we can also put constraints and probe regions of the $n_{s}-r$ parameter space that are within the reach of the current and future CMB experiments. As an example the top right plot in Fig. 14 which shows the complementarity for global cosmic strings at the LISA, constrains the spectral index $n_{s}$ in the range $0.960\lesssim n_{s}\lesssim 0.963$ for the Polynomial Attractor model with $n=1$. A part of the aforementioned range of spectral index (corresponding to $N_{CS}\sim 30-37$) is consistent with Planck and lies well within the reach of next generation CMB experiments like CMB-S4 and LiteBIRD.